# Causality-Respecting Adaptive Refinement for PINNs: Enabling Precise Interface Evolution in Phase Field Modeling

### v1

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a powerful tool for solving physical systems described by partial differential equations (PDEs). However, their accuracy in dynamical systems, particularly those involving sharp moving boundaries with complex initial morphologies, remains a challenge. This study introduces an approach combining residual-based adaptive refinement (RBAR) with causality-informed training to enhance the performance of PINNs in solving spatio-temporal PDEs. Our method employs a three-step iterative process: initial causality-based training, RBAR-guided domain refinement, and subsequent causality training on the refined mesh. Applied to the Allen-Cahn equation, a widely-used model in phase field simulations, our approach demonstrates significant improvements in solution accuracy and computational efficiency over traditional PINNs. Notably, we observe an ’overshoot and relocate’ phenomenon in dynamic cases with complex morphologies, showcasing the method’s adaptive error correction capabilities. This synergistic interaction between RBAR and causality training enables accurate capture of interface evolution, even in challenging scenarios where traditional PINNs fail. Our framework not only resolves the limitations of uniform refinement strategies but also provides a generalizable methodology for solving a broad range of spatio-temporal PDEs. The simplicity and effectiveness of our RBAR-causality combined PINN offer promising potential for applications across various physical systems characterized by complex, evolving interfaces.
@@ 1 Introduction
Within the broader category of DL techniques, physics-informed neural networks (PINNs) [1, 2] have gained particular attention as an alternative to traditional PDE solvers. PINNs approximate PDE solutions by training a deep neural network (DNN) to minimize a loss function that incorporates initial and boundary conditions, as well as the PDE residual at selected points (collocation points). This approach is essentially a mesh-free technique that converts the problem of directly solving governing equations into a loss function optimization problem. PINNs integrate the mathematical model into the network without requiring labeled data, reinforcing the loss function with a residual term from the governing equation that acts as a penalizing term, thus ensuring the network considers the underlying physics rather than solely fitting data. Numerous variants of PINNs have been developed including conservative PINNs [3], sparse PINNs [4], varitaional PINNs [5], domain-decomposition PINNs (XPINNs) [6], and separable PINNs [7], stochastic PINNs [8] to name a few. However, PINNs are not the only methods that constrain ML models with physics. Other popular methods that use DNNs to solve PDEs include the Deep Ritz method [9] and approaches based on the Galerkin or Petrov-Galerkin method [10] including the Deep Galerkin Method [11]. When a Galerkin approach is used on collocation points, the framework is a variant of PINNs, i.e. hp-VPINNs [10]. In the Deep Energy Method [2] the loss function minimizes the variational energy of the system at quadrature points. Other methods have used GPs to solve PDEs [12, 13] or have imposed physical constraints on GPs [14].
@@
While PINNs have demonstrated success across various domains, their application to phase field equations has been limited. This scarcity in literature stems from the challenges posed by sharp transition interface in phase field solutions, which evolve over time [15]. Consequently, certain aspects of phase field model solutions prove more challenging to learn, both spatially and temporally. To address these spatio-temporal challenges, researchers have recently implemented sequential training strategies [16, 17]. The work in [18] has shown that these strategies are rooted in the principle of causality, suggests that temporal evolution in dynamic systems may lead PINNs to converge towards erroneous solutions [17]. While a straightforward causality-based training method can help PINNs avoid these erroneous solutions, their predictive capability in forecasting sharp moving layers remains insufficient [15]. Accurate capture of complex morphology in moving layers necessitates adaptive sampling. Residual-based adaptive refinement (RBAR) methods have been proposed, adding points in areas with large residual losses following different sampling rules [19, 20]. However, these refinement methods treat spatial and temporal dimensions equivalently, disregarding the significance of causality.
@@
In this study, we propose a combined RBAR-causality method to solve the Allen-Cahn equations with high accuracy. To our knowledge, this is the first integration of the causality training method and residual-based adaptive refinement within PINNs to address the Allen-Cahn equations. Our specific contributions are:
@@
  • We demonstrate that our straightforward approach can effectively resolve the Allen-Cahn equation, characterized by a complex sharp-moving interface. This eliminates the need for repetitive training cycles or time-step segmentation of the network, as seen in previous studies [15, 21]. The simplicity of our proposed RBAR-causality method facilitates its potential generalization across various spatiotemporal systems.
@@
  • We identify and discuss the ’overshoot and relocate’ phenomenon observed in the RBAR-causality combined PINN. This phenomenon underscores the exceptional collaborative efficiency of RBAR and causality, highlighting the adaptive error correction capability of the combined approach.
@@ 2 Enhanced PINNs: Integrating Residual-Based Adaptive Refinement and Causality Training
where $\Omega\subset\mathbb{R}^{d}$ denotes a bounded domain on which the PDE is defined, $\partial\Omega$ denotes the domain boundary, $u:\Omega\times[0,T]\to\mathbb{R}$ is the desired solution, $\bm{x}\in\Omega$ is a spatial vector variable, $t$ is time, and $\mathcal{N}$ and $\mathcal{B}$ are spatio-temporal differential operators. In Equation 3, $f:\Omega\to R$ is the driving function, $g:\partial{\Omega}\to R$ is the boundary condition function, and $h:\Omega\backslash\partial{\Omega}\to R$ denotes the initial condition. Known as universal function approximators, the DNN can be employed to approximate the solution $u(\bm{x},t;\bm{\theta})$, where $\bm{\theta}$ denotes all tunable parameters (i.e., weights and bias) of the DNN.
To satisfy the conditions expressed in Eqn.(1)–(3), the PINNs model is trained by minimizing the following composite loss function, $\mathcal{L}(\bm{\theta,\lambda_{b},\lambda_{i}})$, defined as:
$\mathcal{L}(\bm{\theta,\lambda_{b},\lambda_{i}})=\mathcal{L}_{r}(\bm{\theta})+\bm{\lambda}_{b}\mathcal{L}_{b}(\bm{\theta})+\bm{\lambda}_{i}\mathcal{L}_{i}(\bm{\theta}),$ (4)
where $\mathcal{L}_{r}$, $\mathcal{L}_{b}$, and $\mathcal{L}_{i}$ are loss terms derived from the PDE itself (Eq. (1)), the boundary condition (Eq. (2)), and the initial condition (Eq.(3)), respectively. The hyperparameters $\lambda_{b}$ and $\lambda_{i}$ provide flexibility in assigning different weights to each term, facilitating their interaction during the training process. These weights can be either user-specified or automatically tuned during training [22, 23]. In this work, we determine the values of these weights using the self-adaptive method [24]. The individual loss terms are defined as follows:
@@ 2.1 Causality training method
To mitigate the issue of the erroneous convergence of PINN caused by the temporal evolution in spatio-temporal equations, the causality training method is proposed by introducing a simple re-formulation of residual loss function [18]. A new weighted residual loss is defined as:
@@
where $t$ denotes time, and the total time is divided into $N_{t}$ intervals, and the residual loss of each interval is weighted by $\omega_{k}$, expressed as:
$\omega_{k}=exp\left[-\epsilon\sum_{m=1}^{k-1}L_{r}(t_{m},\bm{\theta})\right],\text{ for }k=2,3,...N_{t}.$ (9)
The causality parameter $\epsilon$ controls the steepness of the weights $\omega_{k}$. These weights are inversely and exponentially proportional to the cumulative residual loss from previous time steps. As a result, $L_{r}(t_{k},\bm{\theta})$ will not be minimized unless all preceding residual losses ${L_{r}(t_{m},\bm{\theta})}_{m=1}^{k-1}$ are reduced below a predetermined critical threshold. This reformulation ensures that the PINN is well-trained at every time step, effectively incorporating the principle of causality into the training process.
@@ 2.2 Residual-based adaptive refinement
To achieve a more accurate solution, we propose a Residual-Based Adaptive Refinement (RBAR) method that dynamically refines areas with large residual loss at each time step. Figure 1 illustrates the steps in our framework. Our adaptive h-refinement scheme, which discretizes the problem domain into multiple elements to achieve higher accuracy, is inspired by Goswami et al. [25], while the residual-based refinement approach is influenced by Wu et al. [20]. In the context of causality, we apply refinement exclusively to the spatial dimension. The RBAR method identifies areas of the computational domain with the largest residual loss at each time step. This refinement typically occurs at sharp interfaces to enhance optimization. The RBAR scheme comprises three main steps:
@@
    • Partition the computational domain into elements, each composed of four adjacent points.
    • Calculate the element’s residual loss as the sum of its constituent points’ losses.
@@
    • Sort elements in descending order of their residual loss.
    • Mark the top $4\rho j\%$ of initial points, which account for the most significant residual error, for refinement.
    • Ensure elements do not exceed a predefined maximum refinement level ($R_{\text{max}}$).
    • Record the refinement level of each marked element.
@@
    • Apply h-refinement to marked elements.
    • Add new points at the midpoint of each edge and at the center of the marked element.
    • Replace the original element with four new elements.
    • Repeat this refinement process multiple times across the computational domain.
Figure 1: Flowchart of the proposed RBAR-causality approach.
@@
Input: $\Omega$: Initial computational domain
$T=\{t_{1},\ldots,t_{N}\}$: Set of time steps
@@
$\rho$: Refinement threshold parameter
$\varepsilon$: Convergence tolerance
Output: $\mathcal{M}$: Optimized PINN model, $\Omega_{f}$: Final refined mesh
1
2Initialize PINN model $\mathcal{M}$ and set $\Omega_{0}=\Omega$;
3 $\mathcal{L}_{0}\leftarrow$ InitialTraining($\mathcal{M},\Omega_{0}$);
4
5for $k\leftarrow 1$ to $M_{r}$ do
@@ 6       $\Omega_{k}\leftarrow\Omega_{k-1}$;
9                   $\mathcal{L}_{r}\leftarrow$ ComputeResidualLoss($\mathcal{M},\Omega_{k},t$);
10                   $S_{j}\leftarrow$ IdentifyRefinementSet($\Omega_{k},\mathcal{L}_{r},\rho,j$);
11                   $\Omega_{k}\leftarrow$ ApplyHRefinement($\Omega_{k},S_{j}$);
12
@@ 11                   $\Omega_{k}\leftarrow$ ApplyHRefinement($\Omega_{k},S_{j}$);
14
@@
16      $\mathcal{M},\mathcal{L}_{k}\leftarrow$ CausalityTraining($\mathcal{M},\Omega_{k}$);
17
18 end for
19
[…]
@@ 16      $\mathcal{M},\mathcal{L}_{k}\leftarrow$ CausalityTraining($\mathcal{M},\Omega_{k}$);
22
23
24 Function InitialTraining($\mathcal{M},\Omega$):
25       Perform initial causality training on $\mathcal{M}$ using $\Omega$;
26       return Initial loss $\mathcal{L}_{0}$;
27
28
29
30Function ComputeResidualLoss($\mathcal{M},\Omega,t$):
31       Calculate residual loss for all points in $\Omega$ at time $t$;
32       return Residual loss $\mathcal{L}_{r}$;
33
34
35
36Function IdentifyRefinementSet($\Omega,\mathcal{L}_{r},\rho,j,R_{max}$):
37       Select top $(4\rho j)\%$ of points with highest $\mathcal{L}_{r}$;
38       Retrieve the refinement level R for every point;
39       Note: R = R+1;
40       if $R<R_{max}$ then
41             refine the corresponding points;
42       end if
43
44      return Refinement set $S_{j}$;
45
46
47
48Function ApplyHRefinement($\Omega,S$):
49       Apply $h$-refinement to points in set $S$;
50       return Updated mesh $\Omega$;
51
52
53
54Function CausalityTraining($\mathcal{M},\Omega$):
55       Perform causality training on $\mathcal{M}$ using refined mesh $\Omega$;
56       return Updated model $\mathcal{M}$, New loss $\mathcal{L}$;
57
58
59
Algorithm 1 Enhanced RBAR-Causality Method
In each causality training cycle, we implement an adaptive learning rate scheduler based on the Stochastic Gradient Descent with Warm Restarts method proposed by Loshchilov and Hutter [26]. This approach dynamically adjusts the learning rate throughout the training process, following a cyclical pattern where the rate initially starts high, gradually decreases to a predefined value or zero, and then restarts. This cyclical strategy offers several advantages: it facilitates efficient exploration of the loss landscape in the early stages of each cycle, allows for fine-tuning and convergence to optimal value as the rate decreases, and helps the model escape suboptimal local minima through periodic restarts. By employing this technique, we significantly enhance the robustness of our RBAR-causality based PINNs framework, effectively mitigating the risk of convergence to erroneous solutions that may arise from consistently high learning rates, while also avoiding the premature convergence often associated with monotonically decreasing rates. Furthermore, this approach aligns well with the iterative nature of our RBAR-Causality method, as each refinement stage can benefit from a fresh cycle of the learning rate schedule, potentially leading to improved adaptation to the refined mesh structure.

### v2

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a powerful tool for solving physical systems described by partial differential equations (PDEs). However, their accuracy in dynamical systems, particularly those involving sharp moving boundaries with complex initial morphologies, remains a challenge. This study introduces an approach combining residual-based adaptive refinement (RBAR) with causality-informed training to enhance the performance of PINNs in solving spatio-temporal PDEs. Our method employs a three-step iterative process: initial causality-based training, RBAR-guided domain refinement, and subsequent causality training on the refined mesh. Applied to the Allen-Cahn equation, a widely-used model in phase field simulations, our approach demonstrates significant improvements in solution accuracy and computational efficiency over traditional PINNs. Notably, we observe an ‘overshoot and relocate’ phenomenon in dynamic cases with complex morphologies, showcasing the method’s adaptive error correction capabilities. This synergistic interaction between RBAR and causality training enables accurate capture of interface evolution, even in challenging scenarios where traditional PINNs fail. Our framework not only resolves the limitations of uniform refinement strategies but also provides a generalizable methodology for solving a broad range of spatio-temporal PDEs. The enhanced performance of the RBAR–causality combined framework demonstrates its strong potential for advancing PINN-based modeling of physical systems characterized by complex, evolving interfaces.
@@ 1 Introduction
Within the broader category of DL techniques, physics-informed neural networks (PINNs) Raissi et al. [2019], Samaniego et al. [2020] have gained particular attention as an alternative to traditional PDE solvers. PINNs approximate PDE solutions by training a deep neural network (DNN) to minimize a loss function that incorporates initial and boundary conditions, as well as the PDE residual at selected points (collocation points). This approach is essentially a mesh-free technique that converts the problem of directly solving governing equations into a loss function optimization problem. PINNs integrate the mathematical model into the network without requiring labeled data, reinforcing the loss function with a residual term from the governing equation that acts as a penalizing term, thus ensuring the network considers the underlying physics rather than solely fitting data. Numerous variants of PINNs have been developed including conservative PINNs Jagtap et al. [2020], sparse PINNs Ramabathiran and Ramachandran [2021], variational PINNs Goswami et al. [2020a], domain-decomposition PINNs (XPINNs) Jagtap and Karniadakis [2020], and separable PINNs Cho et al. [2022], stochastic PINNs Karumuri et al. [2020] to name a few. However, PINNs are not the only methods that constrain ML models with physics. Other popular methods that use DNNs to solve PDEs include the Deep Ritz method Yu and others [2018] and approaches based on the Galerkin or Petrov-Galerkin method Kharazmi et al. [2019] including the Deep Galerkin Method Sirignano and Spiliopoulos [2018]. When a Galerkin approach is used on collocation points, the framework is a variant of PINNs, i.e. hp-VPINNs Kharazmi et al. [2019]. In the Deep Energy Method Samaniego et al. [2020] the loss function minimizes the variational energy of the system at quadrature points. Other methods have used Gaussian Processes (GPs) to solve PDEs Yang and Perdikaris [2018], Pang and Karniadakis [2020] or have imposed physical constraints on GPs Swiler et al. [2020].
@@
While PINNs have demonstrated success across various domains, their application to phase field equations has been limited. This scarcity in literature stems from the challenges posed by sharp transition interface in phase field solutions, which evolve over time Wight and Zhao [2020]. Consequently, certain aspects of phase field model solutions prove more challenging to learn, both spatially and temporally. To address these spatio-temporal challenges, researchers have recently implemented sequential training strategies Krishnapriyan et al. [2021], Mattey and Ghosh [2022]. The work in Wang et al. [2022a] has shown that these strategies are rooted in the principle of causality, suggests that temporal evolution in dynamic systems may lead PINNs to converge towards erroneous solutions Mattey and Ghosh [2022]. While a straightforward causality-based training method can help PINNs avoid these erroneous solutions, their predictive capability in forecasting sharp moving layers remains insufficient Wight and Zhao [2020]. Accurate capture of complex morphology in moving layers necessitates adaptive sampling of residual points over the computational domain. Residual-based adaptive refinement (RBAR) methods have been proposed, adding points in areas with large residual losses following different sampling rules Yu et al. [2022], Wu et al. [2023]. However, these refinement methods treat spatial and temporal dimensions equivalently, disregarding the significance of causality.
@@
In this study, we propose a combined RBAR-causality method to solve the Allen-Cahn equations with high accuracy. To the best of our knowledge, this work presents the first integration of causality-based training and residual-based adaptive refinement within physics-informed neural networks (PINNs) to achieve accurate surrogate modeling of solutions with sharp interfaces. We demonstrate the effectiveness of the proposed approach using the Allen–Cahn equation. Our specific contributions are:
@@
  • We demonstrate that our straightforward approach can effectively resolve the Allen-Cahn equation, characterized by a complex sharp-moving interface. This eliminates the need for repetitive training cycles or time-step segmentation of the network, as seen in previous studies Wight and Zhao [2020], Jung et al. [2024]. The simplicity of our proposed RBAR-causality method facilitates its potential generalization across various spatiotemporal systems.
@@
  • We identify and discuss the ‘overshoot and relocate’ phenomenon observed in the RBAR-causality combined PINN. This phenomenon underscores the exceptional collaborative efficiency of RBAR and causality, highlighting the adaptive error correction capability of the combined approach.
@@ 2 Enhanced PINNs: Integrating Residual-Based Adaptive Refinement and Causality Training
where $\Omega\subset\mathbb{R}^{d}$ denotes a bounded domain on which the PDE is defined, $\partial\Omega$ denotes the domain boundary, $u:\Omega\times[0,T]\to\mathbb{R}$ is the desired solution, $\bm{x}\in\Omega$ is a spatial vector variable, $t$ is time, and $\mathcal{N}$ and $\mathcal{B}$ are spatio-temporal differential operators. In Equation 3, $f:\Omega\to R$ is the driving function, $g:\partial{\Omega}\to R$ is the boundary condition function, and $h:\Omega\backslash\partial{\Omega}\to R$ denotes the initial condition. In PINN, the assumed parametric form of the solution is constructed using DNN as a global space-time function approximator. Specifically, the solution $u(\bm{x},t)$ is approximated as:
$u(\bm{x},t)\approx u_{NN}(\bm{x},t;\bm{\theta}).$ (4)
where $\bm{\theta}$ denotes all trainable network parameters that globally define the solution over the entire space–time domain. The neural network representing $u_{NN}(\bm{x},t;\bm{\theta})$ is a fully connected feed-forward network with five hidden layers, each containing 100 neurons and using the hyperbolic tangent activation function. The network weights are initialized using the Xavier Normal method.
To satisfy the conditions expressed in Eqn.(1)–(3), the PINNs model is trained by minimizing the following composite loss function, $\mathcal{L}(\bm{\theta},\beta_{b},\beta_{i})$, defined as:
$\mathcal{L}(\bm{\theta},\beta_{b},\beta_{i})=\mathcal{L}_{r}(\bm{\theta})+\beta_{b}\mathcal{L}_{b}(\bm{\theta})+\beta_{i}\mathcal{L}_{i}(\bm{\theta}),$ (5)
where $\mathcal{L}_{r}$, $\mathcal{L}_{b}$, and $\mathcal{L}_{i}$ are loss terms derived from the PDE itself (Eq. (1)), the boundary condition (Eq. (2)), and the initial condition (Eq.(3)), respectively. The hyperparameters $\beta_{b}$ and $\beta_{i}$ provide flexibility in assigning different weights to each term, facilitating their interaction during the training process. These weights can be either user-specified or automatically tuned during training Wang et al. [2021, 2022b]. In this work, we determine the values of these weights using the self-adaptive method McClenny and Braga-Neto [2020]. The individual loss terms are defined as follows:
@@
In this work, the $L^{2}$-norm (mean squared error) is adopted for all the loss functions $\mathcal{L}_{*}$ ($\in\{r,b,i\}$) . This choice is motivated not only by its differentiability and optimization-friendly properties but also by its physical interpretation: minimizing the $L^{2}$-norm of the PDE residual is analogous to minimizing the ‘energy’ of the violation of the physical laws across the computational domain. This strongly penalizes large deviations from the governing equations, thereby enforcing the physical constraints more effectively.
@@ 2.1 Causality training method
To mitigate the issue of the erroneous convergence of PINN caused by the temporal evolution in spatio-temporal equations, the causality training method is proposed by defining temporal residual loss function $\mathcal{L}_{r}(t,\bm{\theta})$ with spatial set $\left\{\bm{x}^{j}_{r}\right\}_{j=1}^{N_{x}}$ Wang et al. [2022a]:
$\mathcal{L}_{r}(t,\bm{\theta})=\frac{1}{N_{x}}\sum_{j=1}^{N_{x}}(\mathcal{N}[u(\bm{x}_{r}^{j},t;\bm{\theta})]-f(\bm{x}_{r}^{j},t))^{2},$ (9)
where $t$ denotes time, and the total time is divided into $N_{t}$ intervals ($\left\{t_{k}\right\}_{k=1}^{N_{t}}$). Thus, the reformulation of the residual loss function is given by:
@@
the residual loss of each interval is weighted by $\omega_{k}$, expressed as:
$\omega_{k}=exp\left[-\epsilon\sum_{m=1}^{k-1}\mathcal{L}_{r}(t_{m},\bm{\theta})\right],\text{ for }k=2,3,...N_{t}.$ (11)
The causality parameter, $\epsilon=10$, in this work, controls the steepness of the weights $\omega_{k}$. These weights are inversely and exponentially proportional to the cumulative residual loss from previous time steps. As a result, $\mathcal{L}_{r}(t_{k},\bm{\theta})$ will not be minimized unless all preceding residual losses ${\mathcal{L}_{r}(t_{m},\bm{\theta})}_{m=1}^{k-1}$ are reduced below a predetermined critical threshold. This reformulation ensures that the PINN is well-trained at every time step, effectively incorporating the principle of causality into the training process.
@@ 2.2 Residual-based adaptive refinement
To achieve a more accurate solution, we propose a Residual-Based Adaptive Refinement (RBAR) method that dynamically refines areas with large residual loss at each time step. Figure 1 illustrates the steps in our framework. Our adaptive h-refinement scheme, which discretizes the problem domain into multiple elements to achieve higher accuracy, is inspired by Goswami et al. Goswami et al. [2020b], while the residual-based refinement approach is influenced by Wu et al. Wu et al. [2023]. In this mesh-free context, the domain is partitioned into ‘elements’ solely for the purpose of localized residual evaluation and ranking. To simplify the description, this conceptual region is denoted as ‘element’ the in methodology section. In the context of causality, we apply refinement exclusively to the spatial dimension. The RBAR method identifies areas of the computational domain with the largest residual loss at each time step. This refinement typically occurs at sharp interfaces to enhance optimization. The RBAR scheme comprises three main steps:
@@
    • Partition the computational domain into ‘elements’ , each composed of four adjacent points.
    • Calculate the residual loss of element as the sum of its constituent points’ losses.
@@
    • Sort ‘elements’ in descending order of their residual loss.
    • At each adaptive refinement cycle $j$, the top $4\rho j\%$ of the originally discretized elements, corresponding to the largest residual errors, are marked for refinement, where $\rho=5$ is a fixed ratio parameter and $j\in\{1,2,\ldots,M_{a}\}$ with $M_{a}=5$ denoting the maximum number of adaptive refinements per time step.
    • Ensure ‘elements’ do not exceed a predefined maximum refinement level ($R_{\text{max}}$).
    • Record the refinement level of each marked ‘element’.
@@
    • Apply h-refinement to marked ‘elements’.
    • Add new points at the midpoint of each edge and at the center of the marked ‘element’.
    • Replace the original ‘element’ with four new ‘elements’.
    • Repeat this refinement process 5 times across the computational domain.
Figure 1: Conceptual workflow of a mesh-free RBAR–PINN framework using causality training and residual-based adaptive refinement.
@@
Input: $\Omega$: Initial computational domain (collocation elements)
$T=\{t_{1},\ldots,t_{N_{t}}\}$: Discrete time instances
@@
$\rho$: Refinement ratio parameter
$R_{\max}$: Maximum refinement level
Output: $\mathcal{M}$: Trained PINN model, $\Omega_{f}$: Final refined domain
1 Initialize PINN model $\mathcal{M}$ with parameters $\bm{\theta}$;
2 Set initial domain $\Omega_{0}\leftarrow\Omega$;
3 Initial training:
4 Train $\mathcal{M}$ on $\Omega_{0}$ using causality-based loss minimization;
5 for $k\leftarrow 1$ to $M_{r}$ do
@@ 6 $\Omega_{k}\leftarrow\Omega_{k-1}$;
9         Compute residual loss $\mathcal{L}_{r}(\bm{x},t)$ for all elements in $\Omega_{k}$;
10         Sort elements in descending order of $\mathcal{L}_{r}$;
11         Mark the top $4\rho j\%$ of the originally discretized elements for refinement;
12       foreach marked element do
13            Increase refinement level $R\leftarrow R+1$;
14          if $R\leq R_{\max}$ then
15               Apply $h$-refinement to the element;
16
17            end if
18
19         end foreach
20
@@ 15               Apply $h$-refinement to the element;
22
23   end foreach
24 Causality training:
25   Retrain $\mathcal{M}$ on the refined domain $\Omega_{k}$ using causality-weighted temporal loss;
26
@@ 25   Retrain $\mathcal{M}$ on the refined domain $\Omega_{k}$ using causality-weighted temporal loss;
28Set $\Omega_{f}\leftarrow\Omega_{k}$;
@@
Algorithm 1 RBAR–Causality Training Procedure
In each causality training cycle, we implement an adaptive learning rate scheduler based on the Stochastic Gradient Descent with Warm Restarts method proposed by Loshchilov and Hutter Loshchilov and Hutter [2016]. The learning rate, $l_{r}$, is dynamically adjusted using a cosine decay schedule: it starts at $l_{r}=5\times 10^{-3}$, gradually decreases to a predefined minimum or zero, and then restarts in successive cycles to facilitate stable training. This cyclical strategy offers several advantages: it facilitates efficient exploration of the loss landscape in the early stages of each cycle, allows for fine-tuning and convergence to optimal value as the rate decreases, and helps the model escape suboptimal local minima through periodic restarts. By employing this technique, we significantly enhance the robustness of our RBAR-causality based PINNs framework, effectively mitigating the risk of convergence to erroneous solutions that may arise from consistently high learning rates, while also avoiding the premature convergence often associated with monotonically decreasing rates. Furthermore, this approach aligns well with the iterative nature of our RBAR-Causality method, as each refinement stage can benefit from a fresh cycle of the learning rate schedule, potentially leading to improved adaptation to the refined mesh structure.